OpenStudy (anonymous):

Find the solution of \[y'=5 y(3-y)\] with \[y(0)=9\].

OpenStudy (anonymous):

\[y'=5 y(3-y)\] with \[y(0)=9\]

OpenStudy (anonymous):

so, how to antidifferentiate \[1/(3y-y^2)\]

OpenStudy (zarkon):

partial fractions

OpenStudy (zarkon):

\[\frac{1}{y(3-y)}=\frac{1}{3y}+\frac{1}{3(3-y)}\]

OpenStudy (anonymous):

aha! partial fractions? that was fast, dude.

OpenStudy (anonymous):

so \[\int\limits[1/3y+1/(3(3−y))] = \ln(3y)/3 - \ln(9-3y)/3\] is what i got

OpenStudy (zarkon):

\frac{}{}

OpenStudy (anonymous):

hey how do you get the eq editor to show top/bottom fractions like that?

OpenStudy (anonymous):

oh cool, ty.

OpenStudy (anonymous):

could i say \[ln(3y) - ln(9-3y) = ln(3y/(9-3y)\]?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

you can make it look nice by doing this \[\ln\left(\frac{3y}{9-3y}\right)\] \ln\left(\frac{3y}{9-3y}\right)

OpenStudy (anonymous):

so i got \[\frac{y}{3-y} = -\frac{3}{2}*e^{15t}\]

OpenStudy (anonymous):

looks like it will be implicit

OpenStudy (anonymous):

that cant be right, it requires an answer in the form of y(t) = ...

OpenStudy (anonymous):

Here is a different method. the answer you got above can be solved for y, and i think it will produce what I got. Im checking it out right now.

OpenStudy (anonymous):

yeah, what you have above my solution produces the same solution.

OpenStudy (zarkon):

\[\frac{y}{3-y} = -\frac{3}{2}*e^{15t}\] \[\frac{y}{3-y} = a\] \[y=(3-y)a\] \[y=3a-ya\] \[y+ya=3a\] \[y(1+a)=3a\] \[y=\frac{3a}{1+a}\]

OpenStudy (anonymous):

from there, divide the numerator and denominator by 3a, and you'll get what i got.

OpenStudy (anonymous):

WOOOOOWWWW!!!! where do you guys come up with this?!

OpenStudy (anonymous):

black magic lol. Diff equations just seems to be a bunch of rules you have to memorize -.- "if it looks like this, do this" "if it looks like that, do that"

OpenStudy (anonymous):

lol. On a related note, Im surprised the prof expects me to reach that conclusion. He never taught us at least half a dozen tricks you two just used.

myininaya (myininaya):

very cute! i wish i could have stayed up for this problem i enjoy SOME differential problems